# Polynomial-like dynamics of analytic maps

### v1

@@ Abstract.
The theory of polynomial-like maps is of fundamental importance in holomorphic dynamics. We study dynamical properties of a larger class of maps. Our main result is that, under some natural conditions, a map of this class has a completely invariant compact set if and only if this set is the filled Julia set of a polynomial-like restriction of the map. We also generalize this result to include maps with non-connected domains of definition.
@@ 1. Introduction and main results
1.1. Polynomial-like mappings
The theory of polynomial-like mappings was developed by Douady and Hubbard [2] to explain, in particular, the fractal nature of non-linear phenomena in holomorphic dynamics. In this paper, proper holomorphic maps which are not polynomial-like are considered and we are interested in conditions that guarantee a polynomial-like restriction of such a map of the same degree.
@@ 1.1. Polynomial-like mappings
By the Straightening Theorem [2], a PL map $g:V^{\prime}\to V$ is hybrid equivalent (i.e., conjugate in some neighborhood of $K_{g}$ by a quasiconformal homeomorphism which is conformal a.e. on $K_{g}$) to a polynomial of degree $d$. This implies that the set of repelling periodic points of the map $g:K_{g}\to K_{g}$ is dense in $\partial K_{g}$, moreover, for every $n\in\mathbb{N}$, the map $g^{n}:K_{g}\to K_{g}$ has $d^{n}$ fixed points (counting multiplicity), every component of the interior of $K_{g}$ is (pre)periodic, and the set of limit points of the set $\{g^{-n}(x)\}_{n\geq 0}$ coincide with $\partial K_{g}$, for all $x\in K_{g}$ with the only exception if $g^{-1}(x)=\{x\}$ (in the latter case, $x$ is the only critical point of $g$, it is fixed by $g$ and has the maximal multiplicity $d$).
In the present paper we study similar maps but give up the condition that the domain of definition is compactly contained in the range. Our main result is that if such a map has a completely invariant set $K$ which is a full compact containing all critical points then the map admits a polynomial-like restriction to a neighborhood of $K$ and $K$ is its filled Julia set, see Theorem 1.1 and Corollary 1.1 for complete claims. In Theorem 1.3 we generalize this result to include maps with non-connected domains of definition.
1.2. Assumptions and notations
Throughout the paper - with the only exception Section 1.4 (see Theorem 1.3 and Remark 1.3) and Section 6 (where Theorem 1.3 is proved) - we fix a proper holomorphic map $g:U^{\prime}\to U$ of some degree $d\geq 2$ as follows:
  • $U^{\prime},U$ are simply connected (perhaps, unbounded) domains in $\mathbb{C}$ such that $U^{\prime}\subset U$ and $U^{\prime}\neq U$,
  • let $C_{g}=\{x:g^{\prime}(x)=0\}$ be the set of all critical points of $g$, then all iterates $g^{n}(c)$, $n\geq 0$, are well-defined for any $c\in C_{g}$ and the postcritical set
    $P_{g}=\overline{\{g^{n}(C_{g}):n=0,1,...\}}$
    is a compact subset of $U^{\prime}$,
  • if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$.
Let $U_{-i}=g^{-i}(U)$, $n=0,1,2,$.... In particular, $U_{0}=U$ and $U_{-1}=U^{\prime}$. Then $\{U_{-i}\}_{i=0}^{\infty}$ is a decreasing sequence of (connected) simply connected domains and $P_{g}\subset\cap_{i\geq 0}U_{-i}$.
Other notations.
  • $BO(x)=\{y:g^{n}(y)=x\mbox{ for some }n\geq 0\}$ is the backward orbit of $x\in U$.
  • $O(x)=\{g^{n}(x):n=0,1,2,$...\}, the forward orbit of $x$ assuming all iterates $g^{i}(x)$, $i\geq 0$ are defined, i.e, $x\in\cap_{i\geq 0}U_{-i}$. In particular, $O(x)$ is defined for any point of a (forward) invariant set $E\subset U^{\prime}$, i.e., such that $g(E)\subset E$). In this case, $\omega(x)$ denotes, as usual, the set of limit points of $O(x)$.
  • Grand orbit of $x$ is
    $GO(x)=\{z:g^{m}(z)=g^{n}(z)\mbox{ for some }m,n\geq 0\},$
    assuming that $O(x)$ is well-defined; in particular, $GO(c)$ is defined for each $c\in C_{g}$,
  • if $X\subset\mathbb{C}$ is compact, let $\hat{X}$ (the topological hull of $X$) be the complement to the unbounded component of $\mathbb{C}\setminus X$. Then $\hat{X}$ is a full compact. Note that if $X\subset\Omega$ where $\Omega$ is simply connected, then $\hat{X}\subset\Omega$.
1.3. Main results

### v2

@@ Abstract.
The theory of polynomial-like maps is of fundamental importance in holomorphic dynamics. We study dynamical properties of a larger class of maps. Our main result is that, under some natural assumptions, a map of this class has a completely invariant compact set if and only if this set is the filled Julia set of a polynomial-like restriction of the map. We also generalize this result to include maps with non-connected domains of definition.
@@ 1. Introduction and main results
Polynomial-like maps were introduced by Douady and Hubbard [4] to explain, in particular, the fractal nature of non-linear phenomena in holomorphic dynamics. In this paper proper holomorphic map which are not polynomial-like are considered and we are interested in conditions that guarantee a polynomial-like (PL) restriction of such a map. We prove in Theorem 1.1 that, in the class of BI maps, see Definition 1.1, a PL restriction exists if and only if the map has a completely invariant compact set $K$ (under some natural restrictions). This gives a rather complete description of the dynamics of BI maps on such a compact and its neighborhood. It would be interesting to study dynamical properties of locally analytic self-maps of a compact set of degree at least two for other classes of maps.
Let
(1.1) $g:U_{-1}\to U_{0}$
be a proper holomorphic map of some degree $d\geq 2$, where $U_{-1},U_{0}$ are Jordan domains. We study such a map under an extra assumption that it is backward invariant, i.e., all further (following the first one $U_{-1}=g^{-1}(U_{0})$) pullbacks $g^{-i}(U_{0})$, $i=2,3,$..., of $U_{0}$ also exist (BI map, see Definition 1.1).
If $\overline{U_{-1}}\subset U_{0}$, the map (1.1) is polynomial-like and is always BI because all pullbacks exists (see Section 1.1 and Example 1.3 there for details). On the other hand, there are two interesting classes of BI maps which are not, in general, PL maps, see Example 2.1 for the first class and Examples 2.2-2.3 for the second one. If, e.g., in (1.1), $U_{-1}\subset U_{0}$, but $\overline{U_{-1}}\cap\partial U_{0}\neq\emptyset$, and the postcritical set is a subset of $U_{-1}$, the map is BI but not PL. On the other hand, another class of BI maps appears naturally in connection with the renormalization. Namely, a BI map can be associated to any simple renormalization of a quadratic polynomial, see Example 2.2. If, moreover, a quadratic polynomial is infinitely renormalizable and robust but with no unbranched complex bounds, in the sense of McMullen [9], a rescaling subsequence of BI maps along renormalizations converges to a limit BI map, see Example 2.3. In fact, the latter examples inspired the present work.
The set $C_{g}:=\{z\in U_{-1}:g^{\prime}(z)=0\}$ of critical points of the map (1.1) is not empty as $U_{-1},U_{0}$ are simply connected and the degree $d\geq 2$. Let
$P_{g}=\overline{\cup_{n>0}g^{n}(C_{g})}$
be the postcritical set of this map whenever it is well-defined, i.e., assuming $g^{n}(C_{g})\subset U_{-1}$ for all $n\geq 0$.
Definition 1.1.
The map (1.1) is called backward invariant (BI map) if there exist a sequence
$U_{0},U_{-1},...,U_{-n},$...
of Jordan domains and a holomorphic extension $g:\cup_{i=1}^{\infty}U_{-i}\to\cup_{i=0}^{\infty}U_{-i}$ of $g:U_{-1}\to U_{0}$ such that the following hold, for every $i=0,1,$...:
  • $g:U_{-i-1}\to U_{-i}$ is a proper holomorphic map of degree $d$,
  • $P_{g}$ is well-defined and $P_{g}\subset U_{-i}$. Moreover, we assume that if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$.
In the sequel, slightly abusing notation,
(1.2) $g:U^{\prime}\to U,\mbox{ where }U^{\prime}=\cup_{i=1}^{\infty}U_{-i},U=\cup_{i=0}^{\infty}U_{-i},$
always refers to the BI map (unless stated otherwise).
We call a BI map $g:U^{\prime}\to U$ trivial if there is an attracting fixed point $a\in\cap_{i\geq 0}U_{-i}$ such that $g^{n}(z)\to a$ for all $z\in U$.
For every $i\geq 0$ and $n\geq 1$, the map $g^{n}:U_{-i-n}\to U_{-i}$ is a proper map of degree $d^{n}$. By $g_{i}^{-n}:U_{-i}\to U_{-i-n}$ we denote the inverse (multi-valued) map. E.g., $g_{0}^{-n}(U_{0})=U_{-n}$.
Remark 1.2.
We do not assume that $g:U^{\prime}\to U$ is a proper map (of degree $d$). But if it is so (as in Example 2.1), then writing $g^{-n}(x)$ (i.e. dropping $i$ in $g_{i}^{-n}(x)$) is not ambiguous and this would allow us to prove a bit more, see Remark 1.6 and the Complement to Corollary 1.5. However, there is no reason to believe that $g:U^{\prime}\to U$ is a proper map in Examples 2.2-2.3.
The following maps (i)-(iii) which are associated to the BI map (1.2) are BI by themselves: (i) the restricted map $g:\cup_{n=1}^{\infty}g_{0}^{-n}(\Omega)\to\cup_{n=0}^{\infty}g_{0}^{-n}(\Omega)$, for every Jordan domain $\Omega$ such that $P_{g}\subset\Omega\subset U_{0}$, (ii) the shifted map $g:\cup_{i=i_{0}+1}^{\infty}U_{-i}\to\cup_{i=i_{0}}^{\infty}U_{-i}$, for every $i_{0}>0$, and (iii) the iterated map $g^{n}:\cup_{i=1}^{\infty}U_{-ni}\to\cup_{i=0}^{\infty}U_{-ni}$, for every $n>1$. All these maps share the same postcritical set $P_{g}$.
1.1. Polynomial-like maps
@@
The compact $K_{g}$ is connected if and only if $C_{g}\subset K_{g}$ where $C_{g}=\{z\in V^{\prime}:g^{\prime}(z)=0\}$. Equivalently, $P_{g}\subset V$.
By the Straightening Theorem [4], a PL map $g:V^{\prime}\to V$ is hybrid equivalent (i.e., conjugate in some neighborhood of $K_{g}$ by a quasiconformal homeomorphism which is conformal a.e. on $K_{g}$) to a polynomial of degree $d$. Since the polynomial dynamics is well-understood, see e.g., [2], this gives a complete description of the dynamics on $K_{g}$. In particular, the set of repelling periodic points of the map $g:K_{g}\to K_{g}$ is dense in $\partial K_{g}$, for every $n\in\mathbb{N}$, the map $g^{n}:K_{g}\to K_{g}$ has $d^{n}$ fixed points (counting multiplicity), every component of the interior of $K_{g}$ is (pre)periodic [16], etc. Moreover, the set of limit points of $\{g^{-n}(x)\}_{n\geq 0}$ coincide with $\partial K_{g}$, for all $x\in K_{g}$ with the only exception if $g^{-1}(x)=\{x\}$ (in the latter case, $C_{g}=\{x\}$ and is completely invariant).
[…]
Every PL map is, essentially, a BI map:
Example 1.3.
Given a PL map $g:V^{\prime}\to V$ with $P_{g}\subset V^{\prime}$, choose any Jordan domain $U_{0}\subset V$ which is slightly smaller than $V$ (if $V$ is Jordan itself, one can take $U_{0}=V$). Then $U_{-i}:=g^{-i}(U_{0})$, $i=0,1,2,$..., is a well-defined (decreasing in this case) sequence of Jordan domains such that $P_{g}\subset U_{-i}$ and $g:U_{-i-1}\to U_{-i}$ is proper degree $d$, for each $i$. Assume additionally that $C_{g}$ is not a completely invariant singleton (equivalently, $g:V^{\prime}\to V$ is not hybrid equivalent to the power map $z\mapsto z^{d}$). Then $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a BI map.
See Section 2 for examples of BI maps which are not polynomial-like.
Example 1.3 makes natural the following
Definition 1.4.
We say that a BI map $g:U^{\prime}\to U$ has a polynomial-like restriction if there exists a topological disk $V\subset U_{0}$ as follows: if $V^{\prime}:=g_{0}^{-1}(V)$ then $P_{g}\subset V^{\prime}\subset\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a polynomial-like map of degree $d$.
In this case, the filled Julia set $K$ of the PL map $g:V^{\prime}\to V$ is a full completely invariant compact containing $P_{g}$.
1.2. Main result
Example 1.3 provides us a necessary condition for a BI map to have a PL restriction: the existence of a full completely invariant compact set that contains all critical points. We prove that this necessary condition is also sufficient:
@@ 1.2. Main result
Suppose that $g:U_{-1}\to U_{0}$ is a proper holomorphic map of some degree $d\geq 2$ which extends to a BI map $g:U_{-i-1}\to U_{-i}$, $i=1,2,$.... Let $K\subset\cap_{i\geq 0}U_{-i}$ be a compact set as follows:
  1. (t1)
    $K$ is completely invariant,
  2. (t2)
    $K$ is a full compact and $P_{g}\subset K$.
Then $g$ has a polynomial-like restriction and $K$ is its filled Julia set. In particular, $K$ is a continuum,
In [5] we apply Theorem 1.1 to study holomorphic/meromorphic invariant line fields of BI maps. In the present paper we derive the following corollary and its complement.
Given a BI map $g:U^{\prime}\to U$ we define, for each $n\in\mathbb{N}$, the set
(1.3) $R_{n}=\{z|z\in U_{-n}\cap U_{0},g^{n}(z)=z,|(g^{n})^{\prime}(z)|>1\}.$
Note that $R_{n}$ is finite in any compact subset of $U_{0}$, by the Uniqueness theorem. Let
(1.4) $R^{\prime}=\{z|\exists z_{i}\to z,z_{i}\in R_{n_{i}},n_{i}\to\infty\}.$
Corollary 1.5.
For a non-trivial BI map $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$, the following alternative holds:
(a) it does not have a completely invariant full compact $K$ such that $P_{g}\subset K\subset\cap_{i\geq 0}U_{-i}$, and the set $R^{\prime}$ contains a (non trivial) continuum which has common points with $U_{0}$ as well as with its boundary $\partial U_{0}$.
(b) $g$ admits such $K$ and there exists a neighborhood of $W$ of $K$, such that $g:g_{0}^{-1}(W)\to W$ is a PL restriction and $K$ its filled Julia set. Besides,
  1. (1)
    $R^{\prime}\cap K=\partial K$,
  2. (2)
    $R_{n}\cap(W\setminus K)=\emptyset$, for every $n\in\mathbb{N}$,
  3. (3)
    $R^{\prime}\setminus K\subset\partial U_{0}$.
Remark 1.6.
Assume that a BI map $g:U^{\prime}\to U$ is a degree $d$ proper map. Then $\cup_{n\geq 1}R_{n}\subset R^{\prime}$, in particular, in case (b), $g$ has no repelling periodic orbits outside of $K$. Indeed, $g:U_{-i-1}\to U_{-i}$ has degree $d$ and if also $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ has degree $d$ then $g^{-1}(U_{-i})=U_{-i-1}$. This implies that $g_{0}^{-n}(x)=g^{-n}(x)$ for all $x\in U_{0}$ and all $n\in\mathbb{N}$. If now $z\in U_{-n}\cap U_{0}$ is such that $g^{n}(z)=z$, then $g^{ni}(z)=z$, hence, $z\in U_{-ni}$. Therefore, letting $z=z_{i}$ and $n_{i}=ni$ in (1.4) we see that $z\in R^{\prime}$.
This remark is relevant for, e.g., a class of BI maps of Example 2.1. Let us state this explicitly as the following
Complement to Corollary 1.5. Let $V^{\prime},V$ be simply connected domains in $\mathbb{C}$ such that $^{\prime}\subset V, VV$$^{\prime}\neq V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$ and the postcritical set $P_{g}\subset V^{\prime}$. Let $R^{\prime}$ be the closure of all repelling cycles of $g:V^{\prime}\to V$. Then the alternative (a)-(b) holds. Moreover, in case (b) of the alternative, $R^{\prime}\subset K$, and for any $x\in V^{\prime}\setminus K$, either $g^{n}(x)\in V\setminus V^{\prime}$ for some $n\geq 0$, or the forward orbit $O(x)$ of $x$ is well-defined but
$\omega(x)\subset\partial V,$
where $\omega(x)$ is the limit set of $O(x)$.
Next remark is, essentially, a folklore:
Remark 1.7.
For a proper map (1.1), conditions (t1)-(t2) along are not enough for the conclusion of Theorem 1.1 to hold. Indeed, let’s assume, along with (t1)-(t2), that $g:U_{-1}\to U_{0}$ has an attracting or parabolic cycle such that its basin of attraction intersects $U_{-1}\setminus K$. Then it is easy to see that the conclusion of Theorem 1.1 breaks down. Nevertheless, if, additionally, to (t1)-(t2), $K$ is connected and attracting/parabolic cycles as above do not exist, then the conclusion of Theorem 1.1 does hold (see [1] for details, see also Case (B) in the proof of Proposition 3.2 below). The proof of this is heavily based on the fact that $K$ is a connected full compact set. This allows us to conjugate the map $g:U_{-1}\setminus K\to U_{0}\setminus K$ (by uniformazing $\mathbb{C}\setminus K$) to a real analytic covering map on an outer neighborhood of the unit circle $S^{1}$ with no critical points, extend it through $S^{1}$ by the reflection principle and then apply a non-trivial Mane’s theorem which says that a $C^{2}$ covering map of the circle without critical points is expanding away from its attracting and parabolic cycles.
The proof of Theorem 1.1 is a mixture of dynamical and geometric arguments. The main step constitutes the following
Theorem 1.2.
Under the conditions of Theorem 1.1, the set $R^{\prime}$ (see 1.3-1.4) has no points in $U_{0}\setminus K$.
The proof of Theorem 1.2 in the case of connected $K$ is rather simple because we can still use the conjugation of the map $g:U_{-1}\setminus K\to U_{0}\setminus K$ to a circle map (although the proof is more elementary than the one outlined in the above Remark 1.7 as Mane’s theorem is not needed), see Sect. 4.1. Most part of the paper occupies the proof of Theorem 1.2 in the case of disconnected $K$ along with Appendices A-C where we collect auxiliary results which are used in the proof.
1.3. Generalized BI maps
Note that a priori the completely invariant compact $K$ of Theorem 1.1 does not have to be connected although a posteriori it does.
Let us mention the following generalization of Theorem 1.1 where $K$ is never connected. Let, in Definition 1.1, $U_{0}$ be still a Jordan domain while $U_{-1}$ be an open set consisting of finitely many Jordan components $U_{-1}^{k}$, k=1,,mk=1,...,m, for some $m\geq 2$. Let $g:U_{-1}\to U_{0}$ be a proper map of degree $d\geq 2$ meaning that each $g:U_{-1}^{k}\to U_{0}$ is a proper map of some degree $d_{k}\geq 1$ such that $d=\sum_{k}d_{k}$. We say that $g$ extends to a generalized BI map if there exist a sequence $U_{-1},...,U_{-i},$... of bounded open sets each of which consists of finitely many Jordan components, and a holomorphic extension $g:\cup_{i=1}^{\infty}U_{-i}\to\cup_{i=0}^{\infty}U_{-i}$ of $g:U_{-1}\to U_{0}$ such that, for every $i=0,1,$...: (1) $g(U_{-i-1})=U_{-i}$ and, for every component $U$ of $U_{-i}$, $g:U_{-i-1}\to U$ is a proper holomorphic map of degree $d$, (2) either $g$ has no critical points or $P_{g}$ is well-defined, contains at least two points and $P_{g}\subset U_{-i}$.
Recall that a generalized PL map [7] $G:V^{\prime}\to V$ is a holomorphic map between open sets $V^{\prime},V$, where $V$ is simply connected, $\overline{V^{\prime}}\subset V$, and $V^{\prime}$ consists of finitely many simply connected components $V_{k}^{\prime}$, k=1,,Mk=1,...,M, with pairwise disjoint closures, such that each restriction $G:V_{k}^{\prime}\to V$ is a proper map of some degree $D_{k}\geq 1$. Its set of non-escaping points (filled Julia set) $K_{G}=\{z:G^{n}(z)\in V^{\prime},n=0,1,$...\} is a completely invariant full compact. (Note that if $V^{\prime}$ is connected, $G:V^{\prime}\to V$ is the classical (Douady-Hubbard) PL map.) The Straightening theorem also holds: $G:V^{\prime}\to V$ is conjugated, in a neighborhood of $K_{G}$, to a polynomial of degree $D=\sum_{k=1}^{M}D_{k}$ (see e.g. [6] for a proof of a similar claim). If $M\geq 2$, then $K_{G}$ is never connected.
Theorem 1.3.
Suppose that $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a generalized BI map. Let $K\subset\cap_{i\geq 0}U_{-i}$ be a full completely invariant compact set such that either $g$ has no critical points or $P_{g}\subset K$. Then there exist a simply connected domain $V_{0}$, $\overline{V_{0}}\subset U_{0}$, and $m\geq 1$ such that $\overline{V_{-m}}\subset V_{0}$, where $V_{-m}=g_{0}^{-m}(V_{0})$, and $K$ is the filled Julia set of the generalized PL map $g^{m}:U_{-m}\to U_{0}$ of degree $d^{m}$. Furthermore, there exists a neighborhood $V\subset U_{0}$ of $K$ consisting of finitely many pairwise disjoint with their closures simply connected domains such that $\overline{g_{0}^{-1}(V)}\subset V$ and $g:g^{-1}(V)\to V$ is a proper map of degree $d$.
1.4. Description of the paper
Section 2 contains examples of BI maps which are not polynomial-like.
In Section 3 some general properties of BI maps are collected. Primarily, we prove in Proposition 3.2 that any BI map which is not trivial (i.e., not a basin of attraction of an attracting fixed point) either has a lot of repelling cycles close to the boundary and does not admit a PL restriction, or has a PL restriction, with filled Julia set containing almost all repelling cycles.
Theorem 1.2 is proved in Section 4. The proof in the case of disconnected compact $K$ is based on general facts and their corollaries from Appendices A-C.
Short Sections 5-6-7 contain proofs of Theorem 1.1 (as a simple consequence of Theorem 1.2), Corollary 1.5 along with its Complement, and Theorem 1.3, respectively.
Appendix A. We prove following closely [8], [13] that a locally analytic self-map of a compact set $K$ of degree $d\geq 2$ has topological entropy $\geq\log d$. Coupled with the Variational Principle and the formula for the Hausdorff dimension of invariant measures, this implies that the logarithmic capacity of the completely invariant compact set is positive.
Appendix B. Given a (disconnected) compact $X\subset\mathbb{C}$ of positive logarithmic capacity and a universal covering $\pi:\mathbb{D}\to\mathbb{C}\setminus X$, we recall a classical result about the existence and properties of an infinite Blaschke product $B:\mathbb{D}\to\mathbb{D}$ such that its projection is Green’s function of $\mathbb{D}\setminus X$. In Corollary B.2 we use Lusin-Privalov’s construction to build a nice simply connected domain $W\subset\mathbb{D}$, in particular, the length of $\partial W\cap\partial\mathbb{D}$ is positive, $\pi(W)$ is contained in a given neighborhood of $X$ and $\pi(w)\to X$ as $w\to\partial W\cap\partial\mathbb{D}$ in $W$.
Appendix C. We consider a univalent map $f:\mathbb{D}\to\mathbb{D}$ and show, based on an inequality proved in [10], [12], [15] that, if $E,f(E)\subset\partial\mathbb{D}$, then the length of $E$ is small provided $|f^{\prime}(0)|$ is small and $f(0)$ is away from the unit circle.
Acknowledgments. I would like to thank Feliks Przytycki for several fruitful discussions concerning Theorem A.1, and Misha Sodin, Sebastian van Strien and Vladlen Timorin for helpful comments and references.
2. Examples of BI maps
The next example generalises the previous Example 1.3. We obtain BI maps by relaxing some conditions in the definition of polynomial-like maps and in the choice of $U_{0}$.
Example 2.1.
Let $V^{\prime},V$ be simply connected (perhaps, unbounded) domains in $\mathbb{C}$ such that $^{\prime}\subset V, VV$$^{\prime}\neq V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$. We don’t assume that $V^{\prime}$ is compactly contained in $V$. Such a map must have a non-empty set of critical points $C_{g}=\{x:g^{\prime}(x)=0\}$. Assume that the postcritical set $P_{g}=\cup_{n>0}g^{n}(C_{g})$ is well-defined and is a compact subset of $V^{\prime}$, and if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$. Let $U_{0}\subset V$ be an arbitrary Jordan domain such that $P_{g}\subset U_{0}$. Define $U_{-i}=g^{-i}(U)$, $n=0,1,2,$.... Then $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a BI map.
Sub-example: if $V^{\prime},V$ above are bounded and $\overline{V^{\prime}}\subset V$ (i.e., $g:V^{\prime}\to V$ is PL), then it is not difficult to see that the obtained BI map has a PL restriction if and only if $K_{g}\subset U_{0}$, where $K_{g}$ is the filled Julia set of the PL map $g:V^{\prime}\to V$ (the restriction is then $g:g^{-n}(V^{\prime})\to g^{-n}(V)$, for some $n>0$ such that $g^{-n}(V)\subset U_{0}$).
The following examples of BI maps are related to renormalizable quadratic polynomials.
Example 2.2.
Let $f(z)=z^{2}+c$ be a quadratic polynomial. Assume that it admits a simple renormalization of a period $n>1$ (see [9] for definitions). Let $J_{n}$ be the filled Julia set of a PL restriction of $g^{n}$ with a single critical point at 0, and let $P_{n}$ be the postcritical set of this PL map. Note that $P_{n}=J_{n}\cap P_{f}$. Let $\gamma_{n}$ be a Jordan curve in $M:=\hat{\mathbb{C}}\setminus P_{f}$ which is homotopic in $M$ to a loop that separates $J_{n}$ from $P_{f}\setminus J_{n}$. Let $U_{0}$ be a simply connected domain bounded by $\gamma_{n}$. Pulling back $U_{0}$ along the renormalization as in [9], we obtain a domain $U_{-1}\supset P_{n}$ and degree two proper restriction map $f^{n}:U_{-1}\to U_{0}$ with the postcritical set $P_{n}$. Similarly, we pull back $U_{-1}$ getting $U_{-2}$, and so on. As a result, we obtain a BI map $f^{n}:U_{-i-1}\to U_{-i}$, $i=0,1,$..., with the postcritical set $P_{n}$
Example 2.3.
Let $f(z)=z^{2}+c$ be infinitely renormalizable. Assume that $f$ is robust and has no unbranched complex bounds ([9], and see below). Then a limit BI map naturally appears as follows. First, at each period $n$ of simple renormalization we have a BI map obtained as in the previous example, where we start with a geodesic $\gamma_{n}$ of the Riemann surface $M$ which is homotopic to a loop in $M$ around $J_{n}$. We rescale this BI map around zero by a linear map $A_{n}$ in such a way that the rescaling postcritical set $P_{n}^{*}=A_{n}(P_{n})$ becomes of size $1$ and obtain, for each such $n$, a rescaled BI map $g_{n}:U_{-i-1}^{(n)}\to U_{-i}^{(n)}$, $i=0,1,2,$..., with the postcritical set $P_{n}^{*}$, where $g_{n}=A_{n}\circ f^{n}\circ A_{n}^{-1}$ is the rescaling of $f^{n}$.
Now, assume that $L_{-}\leq\ell(\gamma_{n})\leq L_{+}$, for some $0<L_{-}<L_{+}<\infty$ and all $n$ along a subsequence $\mathcal{R}$ of simple renormalizations, where $\ell$ measures the hyperbolic length on $M$.
Such maps exists, see [3].
Following technique of [9] (see particularly Theorem 10.12 there), we can show that, for every $i\geq 0$, the sequence of maps $\{g_{n}:U_{-i-1}^{(n)}\to U_{-i}^{(n)}\}_{n\in\mathcal{R}}$ is pre-compact in the Caratheodory topology. It allows us, inductively for each $i=0,1,$..., to pass to a limit along a subsequence of $n$’s and obtain a limit sequence of proper degree two maps $g:U_{-i-1}\to U_{-i}$, $i=0,1,2,$..., alone with a compact $X$ which is a Hausdorff limit point of the sequence $\{P_{n}^{*}\}$. Moreover, $X\subset\cap_{i\geq 0}U_{-i}$, $0\in X$ and, for every $k=1,2,$..., $g_{n}^{k}(0)\to g^{k}(0)\in X$. This implies that all maps $g:U_{-i-1}\to U_{-i}$ share the same postcritical set which is a subset of $X$. Thus $g:U_{-i-1}\to U_{-i}$ ($i\geq 0$) is a BI map.
3. On BI maps
Lemma 3.1.
Let $Y\subset\cap_{i\geq 0}U_{-i}$ be a compact set. The following conditions are equivalent:
  1. (1)
    $g^{-1}(Y)=Y$, i.e., $Y$ is completely invariant,
  2. (2)
    $g_{i}^{-1}(Y)=Y$ for all $i\geq 0$,
  3. (3)
    $g_{i}^{-1}(Y)=Y$ for some $i\geq 0$ and $g_{j}^{-1}(x)=g_{i}^{-1}(x)$ for any $j$ and any $x\in Y$,
  4. (4)
    $g:Y\to Y$ has degree $d$ (counting multiplicities).
Moreover, if either of (1)-(4) holds, then, given $n>0$, for every neighborhood of $W$ of $Y$ there exists another neighborhood $W^{\prime}$ of $Y$ as follows: for every j=1,,nj=1,...,n, two (multi-valued) maps $g_{0}^{-j}=g_{j-1}^{-1}\circ...\circ g_{0}^{-1}$ and $(g_{0}^{-1})^{j}=g_{0}^{-1}\circ...\circ g_{0}^{-1}$ ($j$ times) are well-defined on $W^{\prime}$, $g_{0}^{-j}=(g_{0}^{-1})^{j}$ on $W^{\prime}$ (as multi-valued maps), and $g_{0}^{-j}(W^{\prime})=(g_{0}^{-1})^{j}(W^{\prime})\subset W$.
Proof.